OpenStudy (anonymous):

The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is 40 in long. What is the length of the other leg?

OpenStudy (anonymous):

base^2+altitude^2=Hypotenuse^2.

OpenStudy (anonymous):

then 41^2=40^2+X^2

OpenStudy (anonymous):

X^2=81 --------->x=9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

pleasure to help..:)